OpenStudy (anonymous):

Activity title: solving problems for uniformly accelerated rectilinear motion learning target: to calculate displacement,velocity,acceleration and time using the four kinematic equations. 1. A french high-velocity train accelerated uniformly from 250km/hr to 300 km/hr as it sped through a distance of 50 km.finds its acceleration.

OpenStudy (anonymous):

its hard

OpenStudy (anonymous):

275ms^-2 i gave u the formula

OpenStudy (anonymous):

what is the formula?

OpenStudy (anonymous):

first tell me is ti right?

OpenStudy (anonymous):

v^2 - u^2 = 2as s is the dist

OpenStudy (anonymous):

some 1 answer like this v² = u² + 2as where v = final velocity u = initial velocity s = distance here v = 300 km/h = 300,000 m/3600 s = 250/3 m/s u = 250 km/h = 250,000 m/3600s = 625/9 m/s s = 50 km = 50000 m

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i feel confused but i remembered the prof on the blackboard write 3600s something from the other question..

OpenStudy (anonymous):

ishaan your formula is correct how can i use that?

OpenStudy (anonymous):

hey kingkos, I have already sent u the full solution!!!!!!

OpenStudy (anonymous):

where?

OpenStudy (anonymous):

u keep posting the same question repeatedly and do not wait for answer....... here it is again.... above equation can be rearranged as follows: v² - u² ------ = a 2 s ( 250 )^2 ( 625 )^2 m^2 (------ ) - (------) ------- ( 3 ) ( 9 ) s^2 ----------------------------------- = a 2 * 50000 m 62500 390625 m^2 ------ - -------- ------ 9 81 s^2 -------------------------- = a 100000 m 562,500 - 390,625 m^2 ------------------ ------ 81 s^2 ------------------------- = a 100,000 m 171,875 1 m^2 ------- * ---------- --------- = a 81 100,000 s^2 * m

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

171,875 m --------- ----- = a 8,100,000 s^2 so a = 0.02 m/s^2

OpenStudy (anonymous):

i will study this your answer is interesting thank you sir

OpenStudy (anonymous):

yehey

OpenStudy (anonymous):

how can i be fast to get the conversion ur so fast i need it when the test come